OpenStudy (anonymous):

A laser beam in air is incident upon ethanol at an angle of incidence of 37 degrees. What is the angle of refraction?

OpenStudy (aaronq):

i think you can use snell's law \(\dfrac{sin\theta_1}{sin\theta_2}=\dfrac{\eta_2}{\eta_1}\)

OpenStudy (anonymous):

I'm aware that I have to use Snell's Law, I'm just not certain on how to apply it.

OpenStudy (aaronq):

okay, you need the refractive index of air, and the refractive index of ethanol Air at STP 1.00029 Ethyl alcohol 1.36 from here http://hyperphysics.phy-astr.gsu.edu/hbase/tables/indrf.html then you just plug stuff in. \(\dfrac{sin(37^o)}{sin\theta_2}=\dfrac{1.36}{ 1.00029 }\) 

OpenStudy (anonymous):

You are fantastic, thank you!

OpenStudy (aaronq):

thanks, no problem !